English teacher wants to test whether the mean reading speed of students is 550 words per minute. A sample of 12 students revealed a sample mean of 540 word per minute with a standard deviation of 5 words per minute. At the 0.05 significance level, is the reading speed from 550 words per minute?


1
Expert's answer
2021-07-15T09:39:19-0400

"\mu = 550 \\

n=12 \\

\bar{x}=540 \\

s=5 \\

H_0: \mu = 550 \\

H_1: \mu ≠ 550"

When n<30 and "\sigma" is unknown, a one sample t-test is used.

"t= \frac{\bar{x}- \mu}{s/ \sqrt{n}} \\

t = \frac{540-550}{5 / \sqrt{12}}=-6.93 \\

df=n-1=12-1=11 \\

α=0.05"

Two-tailed test

"t_{\u03b1\/2,df} = t_{0.025, 11}=2.201"

If test statistics "|t| > t_{\u03b1\/2,n-1}" , you have sufficient evidence to reject H0.

"|t|=6.93 > t_{0.025, 11}=2.201"

Reject H0.

There is enough evidence to conclude that the mean reading speed of students is NOT 550 words per minute at the 0.05 significance level.